# Non-Abelian anyons in a periodically-driven Abelian model

### 06 Feb 2024 v1

@@ II Floquet engineering
III Appendix: Exchange phases
Bogoliubov-de Gennes quasiparticles.
In this section, we discuss how the Floquet-BCS ground state of the Floquet Hamiltonian is constructed and used to compute geometric phases when vortices are displaced. The fermion Floquet Hamiltonian takes the general form
$\hat{H}=-\sum_{j}\mu_{j}n_{j}-\sum_{\langle ij\rangle}\big{(}J_{ij}\hat{f}_{i}^{\dagger}\hat{f}_{j}+\Delta^{*}_{ij}\hat{f}_{i}\hat{f}_{j}+\mathrm{H.c.}\big{)}$ (S36)
@@ III Appendix: Exchange phases
|$\ket{\Phi}=\mathcal{N}\exp\left(\frac{1}{2}\sum_{ij}T_{ij}\hat{f}_{i}^{\dagger}\hat{f}_{j}^{\dagger}\right)\ket{0},$ (S51)
with a Thouless matrix $T_{ij}$, with respect to the chosen reference state |$\ket{0_{\rm tc}}$, which is the fermion vacuum. The normalization constant is $\mathcal{N}=\sqrt{\mathrm{det}(u)}$, with $u$ defined in Eq. (S48). In case |$\ket{\Phi}$ has vanishing overlap with the fermion vacuum, $\langle\Phi|0_{\rm tc}\rangle=0$, a different reference state needs be chosen, as also further discussed below. By imposing $\hat{\alpha}_{n}\ket{\Phi}=0$, the Thouless matrix $T_{ij}$ is found as $T=-(u^{*})^{-1}v^{*}$ [48]. Since, for the toric-code model on a torus, only pairs of fermionic quasiparticles can be created or annihilated, the physical states must have even fermionic parity. The ground state parity is determined by the BdG Hamiltonian and is computed following Ref. [8].
@@
Geometric phase. To compute the geometric phase accumulated by moving a vortex along a given path, we follow the strategy of Refs. [21, 45], adapted here to the Floquet system. We consider controlled vortex movement within the bulk, that never spans the whole lattice, such that no flipping of the eigenvalues of the Wilson loop operators (change of superselection sector) can occur. For each position of the vortices along the path, the fermion Floquet-BCS corresponding to that vortex configuration is computed from the BdG Hamiltonian as explained in the previous section, resulting in a sequence of fermion ground states |$\ket{\Phi_{j}}$. To avoid dependence of the phase from the details of the exchange path, Levin-Wen’s protocol [45] is employed, which requires one to compute the difference $\theta=\theta_{P}-\theta_{P^{\prime}}$ of the phases $\theta_{P}$ accumulated along two $T$-shape paths depicted in Fig. 3(a). The phase along path $P$ is computed as
$\theta_{P}=\mathrm{Arg}\prod_{(i,i+1)\in P}\bra{\Phi_{i+1}}\hat{t}_{i+1,i}\ket{\Phi_{i}}$ (S52)
Here, the operator $\hat{t}_{i,j}$ represents the fermionic part of the $e$-boson tunnelling operators of Eq. (S23) which converts the vortex spatial configuration |$\ket{\vec{n}_{j}^{e}}$, associated with the fermion BCS ground state |$\ket{\Phi_{j}}$, into |$\ket{\vec{n}_{j}^{e}}$, associated with |$\ket{\Phi_{i}}$. In the quasiparticle picture, $\hat{t}_{i,j}$ is given by the string $\hat{\tilde{\mathcal{S}}}_{p}^{(y)}$ of fermion parities of Eq. (S24). Hence, the matrix elements in Eq. (S52) reduce, for horizontal vortex displacement, to computing the overlap between ‘neighbouring’ BCS wavefunctions, while, for vertical displacement, one needs to compute matrix elements of the form
@@
where $\mathrm{pf}(\cdot)$ denotes the Pfaffian. To compute the matrix elements of Eq. (S53), we note that, if |$\ket{\Phi}$ is the quasiparticle vacuum associated with the BdG Hamiltonian $\hat{H}_{\rm BdG}$, then |$\ket{\Phi^{\prime}}=\hat{Q}^{\dagger}\ket{\Phi}$ with $\hat{Q}$ unitary is a quasiparticle vacuum associated with the transformed BdG Hamiltonian
$\hat{H}_{\rm BdG}^{\prime}=\hat{Q}^{\dagger}\hat{H}_{\rm BdG}\hat{Q}\ .$ (S66)
Given the matrix $\hat{U}$ of eigenvectors of $\hat{H}_{\rm BdG}$, the matrix $\hat{U}^{\prime}$ of eigenvectors of $\hat{H}_{\rm BdG}^{\prime}$ is then $\hat{U}^{\prime}=\hat{Q}^{\dagger}\hat{U}$. The latter can then be used to construct the representation of the BCS vacuum |$\ket{\Phi^{\prime}}$ with respect to the reference state |$\ket{\Phi_{1}}$ according to Eq. (S54) and (III) and then be used to compute its overlap with |$\ket{\Phi}$ via Eq. (S65). However, there is an additional caveat. Using the representation (S54) fixes the gauge of |$\ket{\Phi^{\prime}}$ by imposing a real $\langle{\Phi^{\prime}|\Phi_{1}}\rangle$, which is unwanted. Indeed, consistency with the gauge choice made along the path, $\langle{\Phi_{i}|\Phi_{1}}\rangle\in\mathbb{R}$, only requests that each ground state |$\ket{\Phi}$ (and not also $\hat{Q}\ket{\Phi}$) has real overlap with the reference state |$\ket{\Phi_{1}}$. To remedy for this, in the computation of the phase in Eq. (S53) we introduce a fictitious intermediate step to |$\ket{\Phi_{i}}$ to fix the gauge choice, namely
@@
To see that this re-establishes the correct phase, note that |$\ket{\Phi^{\prime}}$ is equal to $\hat{Q}\ket{\Phi}$ up to some phase $\theta$, |$\ket{\Phi^{\prime}}=e^{i\theta}\hat{Q}\ket{\Phi}$. Inserting the latter expression into Eq. (S67), one obtains that the phase $\theta$ cancels out, leaving
@@
If $\hat{Q}$ is Hermitian, as $\hat{\tilde{\mathcal{S}}}_{p}^{(y)}$ of Eq. (S53) is, the diagonal element $\langle\Phi_{i}|\hat{Q}|\Phi_{i}\rangle$ is real, and thus the phase contribution in Eq. (S68) only comes from $\langle\Phi_{j}|P|\Phi_{i}\rangle$ as desired.

### 28 Feb 2024 v2

@@ II.1 Effective Hamiltonian
II.2 Numerical determination of the Floquet Hamiltonian
In order to determine the effective Floquet Hamiltonian numerically, the Hamiltonian $\hat{H}_{\psi}(t)$ of Eq. (S28) is first represented in Bogoliubov-de Gennes (BdG) form,
$\hat{H}_{\psi}(t)=(1/2)(\hat{\bm{f}}^{\dagger},\hat{\bm{f}})\hat{H}_{\rm BdG}(t)(\hat{\bm{f}},\hat{\bm{f}}^{\dagger})^{T},$ (S42)
where $\hat{\bm{f}}=\{\hat{f}_{1},\ldots,\hat{f}_{N}\}$ and $\hat{\bm{f}}^{\dagger}=\{\hat{f}_{1}^{\dagger},\ldots,\hat{f}_{N}^{\dagger}\}$, with $N=L\times L$, which involves a time-periodic BdG Hamiltonian $\hat{H}_{\rm BdG}(t)$. The effective Floquet-BdG Hamiltonian $\hat{H}_{F,\mathrm{BdG}}$ generated by $\hat{H}_{\rm BdG}(t)$ according to Floquet’s decomposition [Eq. (S29)] is determined numerically by diagonalizing the end-of-period propagator $\hat{U}_{\rm BdG}(t_{0}+T,t_{0})$ determined through direct time propagation of the Schrödinger Eq.
As discussed in the main text and detailed in section ‘Vortex Exchange Phases’, the calculation of the exchange phases builds on the assumption that ground states of the effective Floquet Hamiltonian for the driven fermions, corresponding to a given distribution of $e$ quasiparticles, are prepared. In the high-frequency regime considered here, all $\hat{H}_{F,\mathrm{BdG}}^{[t_{0}]}$ for any initial time $t_{0}$ and their ground state coincide to leading order in $1/\omega$ [see Eq. (S31b)]: the Hamiltonians are both well approximated by the time-averaged Hamiltonian of Eq. (S36), which possesses entirely the desired topological properties for the effective fermion bands, and only differ by small corrections due to the weak higher-order terms. This is indeed verified by comparing the vortex exchange phases obtained using ground states of $\hat{H}_{\rm FBdG}^{[t_{0}]}$ for different initial time $t_{0}$, as shown in Fig. S4. From Fig. S4, one can appreciate that, for all system sizes shown, the variation in the phase is within $\sim 0.3\%$.
Figure S4: Vortex exchange phase $R_{\psi}^{ee}$ in the $\psi$ sector, divided by its ideal value $83\pi/$ 8 as a function of the global phase $\varphi_{0}$ of the driving functions, defined by $d_{x}(t)=J+2\Delta\cos(\omega t+\varphi_{0})$ and $d_{y}(t)=J+2\Delta\cos(\omega t+\varphi_{0}+\pi/2)$. The relative phase $\pi/2$ imprints $p+ip$ pairing, as discussed in the main text. The driving amplitude and frequency are set at $J=\Delta=\mu_{\psi}/16$ and $\omega=2\mu_{\psi}-4J=7\mu_{\psi}/4$ in the topological region. The gray dotted line is a guide for the eye at the ideal value 1.
III Vortex exchange phases
In this section, we discuss how the ground state of the Floquet Hamiltonian is constructed and used to compute geometric phases when $e$ vortices are displaced.
Bogoliubov-de Gennes quasiparticles. The fermion Floquet Hamiltonian takes the general form
$\hat{H}=-\sum_{j}\mu_{j}n_{j}-\sum_{ij}\big{(}J_{ij}\hat{f}_{i}^{\dagger}\hat{f}_{j}+\Delta^{*}_{ij}\hat{f}_{i}\hat{f}_{j}+\mathrm{H.c.}\big{)}$ (S43)
@@ III Vortex exchange phases
|$\ket{\Phi}=\mathcal{N}\exp\left(\frac{1}{2}\sum_{ij}T_{ij}\hat{f}_{i}^{\dagger}\hat{f}_{j}^{\dagger}\right)\ket{0_{\rm tc}},$ (S58)
with a Thouless matrix $T_{ij}$, with respect to the chosen reference state |$\ket{0_{\rm tc}}$, which is the fermion vacuum. The normalization constant is $\mathcal{N}=\sqrt{\mathrm{det}(u)}$, with $u$ defined in Eq. (S55). In case |$\ket{\Phi}$ has vanishing overlap with the fermion vacuum, $\langle\Phi|0_{\rm tc}\rangle=0$, a different reference state needs be chosen, as also further discussed below. By imposing $\hat{\alpha}_{n}\ket{\Phi}=0$, the Thouless matrix $T_{ij}$ is found as $T=-(u^{*})^{-1}v^{*}$ [53]. Since, for the toric-code model on a torus, only pairs of fermionic quasiparticles can be created or annihilated, the physical states must have even fermionic parity. The ground state parity is determined by the Floquet-BdG Hamiltonian and is computed following Ref. [8]. Namely, the Floquet Hamiltonian is first expressed in terms of Majorana operators [Eq. (S14)] as $\hat{H}_{F}=i\sum_{p,p^{\prime}}{H}^{(\gamma)}_{F,pp^{\prime}}\hat{\gamma}_{p}\hat{\gamma}_{p^{\prime}}^{\prime}$. The ground state parity is then given by the Pfaffian $\mathrm{Pf}(B)$ of the real skew-symmetric matrix $B$ defined as
$B=Q\begin{pmatrix}0&1&&&\\ -1&0&&&\\ &&\ddots&&\\ &&&0&1\\ &&&-1&0\end{pmatrix}Q^{T},$ (S59)
where $Q$ is the matrix decomposing ${H}^{(\gamma)}_{F}$ in the form
${H}^{(\gamma)}_{F}=Q\begin{pmatrix}0&\varepsilon_{1}&&&\\ -\varepsilon_{1}&0&&&\\ &&\ddots&&\\ &&&0&\varepsilon_{N}\\ &&&-\varepsilon_{N}&0\end{pmatrix}Q^{T}.$ (S60)
@@
Geometric phase. To compute the geometric phase accumulated by moving a vortex along a given path, we follow the strategy of Refs. [24, 50], adapted here to the Floquet system. We consider controlled vortex movement within the bulk, that never spans the whole lattice, such that no flipping of the eigenvalues of the Wilson loop operators (change of superselection sector) can occur. For each position of the vortices along the path, the fermion Floquet-BCS corresponding to that vortex configuration is computed from the BdG Hamiltonian as explained in the previous section, resulting in a sequence of fermion ground states |$\ket{\Phi_{j}}$. To avoid dependence of the phase from the details of the exchange path, Levin-Wen’s protocol [50] is employed, which requires one to compute the difference $\theta=\theta_{P}-\theta_{P^{\prime}}$ of the phases $\theta_{P}$ accumulated along two ‘T-shaped’ paths depicted in Fig. 3(a). The phase along path $P$ is computed as
$\theta_{P}=\mathrm{Arg}\prod_{(i,i+1)\in P}\bra{\Phi_{i+1}}\hat{t}_{i+1,i}\ket{\Phi_{i}}.$ (S61)
Here, the operator $\hat{t}_{i,j}$ represents the fermionic part of the $e$-boson tunnelling operators of Eq. (S26) which converts the vortex spatial configuration |$\ket{\vec{n}_{j}^{e}}$, associated with the fermion BCS ground state |$\ket{\Phi_{j}}$, into |$\ket{\vec{n}_{i}^{e}}$, associated with |$\ket{\Phi_{i}}$. In the quasiparticle picture, $\hat{t}_{i,j}$ is given by the string $\hat{\tilde{\mathcal{S}}}_{p}^{(y)}$ of fermion parities of Eq. (S27). Hence, the matrix elements in Eq. (S61) reduce, for horizontal vortex displacement, to computing the overlap between ‘neighbouring’ BCS wavefunctions, while, for vertical displacement, one needs to compute matrix elements of the form
@@
where $\mathrm{pf}(\cdot)$ denotes the Pfaffian. To compute the matrix elements of Eq. (S62), we note that, if |$\ket{\Phi}$ is the quasiparticle vacuum associated with the BdG Hamiltonian $\hat{H}_{\rm BdG}$, then |$\ket{\Phi^{\prime}}=\hat{Q}\ket{\Phi}$ with $\hat{Q}$ unitary is a quasiparticle vacuum associated with the transformed BdG Hamiltonian
$\hat{H}_{\rm BdG}^{\prime}=\hat{Q}\hat{H}_{\rm BdG}\hat{Q}^{\dagger}\ .$ (S75)
Given the matrix $\hat{U}$ of eigenvectors of $\hat{H}_{\rm BdG}$, the matrix $\hat{U}^{\prime}$ of eigenvectors of $\hat{H}_{\rm BdG}^{\prime}$ is then $\hat{U}^{\prime}=\hat{Q}\hat{U}$. The latter can then be used to construct the representation of the BCS vacuum |$\ket{\Phi^{\prime}}$ with respect to the reference state |$\ket{\Phi_{1}}$, according to Eqs. (S63) and (III), and then be used to compute its overlap with |$\ket{\Phi}$ via Eq. (S74). However, there is an additional caveat. Using the representation (S63) fixes the gauge of |$\ket{\Phi^{\prime}}$ by imposing a real $\langle{\Phi^{\prime}|\Phi_{1}}\rangle$, which is unwanted. Indeed, consistency with the gauge choice made along the path, $\langle{\Phi_{i}|\Phi_{1}}\rangle\in\mathbb{R}$, only requests that each ground state |$\ket{\Phi}$ (and not also $\hat{Q}\ket{\Phi}$) has real overlap with the reference state |$\ket{\Phi_{1}}$. To remedy for this, in the computation of the phase in Eq. (S62) we introduce a fictitious intermediate step to |$\ket{\Phi_{i}^{\prime}}$ to fix the gauge choice, namely
@@
To see that this re-establishes the correct phase, note that |$\ket{\Phi^{\prime}}$ (in the BCS representation enforcing a real $\langle{\Phi^{\prime}|\Phi_{1}}\rangle$) is equal to $\hat{Q}\ket{\Phi}$ up to some phase $\theta$, |$\ket{\Phi^{\prime}}=e^{i\theta}\hat{Q}\ket{\Phi}$. Inserting the latter expression into Eq. (S76), one obtains that the phase $\theta$ cancels out, leaving
@@
If $\hat{Q}$ is Hermitian, as $\hat{\tilde{\mathcal{S}}}_{p}^{(y)}$ of Eq. (S62) is, the diagonal element $\langle\Phi_{i}|\hat{Q}|\Phi_{i}\rangle$ is real, and thus the phase contribution in Eq. (S77) only comes from $\langle\Phi_{j}|\hat{Q}|\Phi_{i}\rangle$ as desired.